(c) [ (2–1x)dx $ 6+3) dx (d)Evaluate the given integrals by graphing the integrands and using well known area formulas
Answer #1

(c) (2-Ja) dx 92) 0) (1) 채 - 2 [글(2+1)x] T 3 = Brms / A= I = 2 (1 (S, 빛) (d) +) 지 (0, 3)) (-6, 0) 5 - 글 (+) 4 - 2(은) Ans, 있